OpenStudy (abmon98):

A stone is projected from a point O on horizontal ground with speed V m s−1 at an angle θ above the horizontal, where sinθ = 3 5 . The stone is at its highest point when it has travelled a horizontal distance of 19.2 m. (i) Find the value of V. [3] After passing through its highest point the stone strikes a vertical wall at a point 4 m above the ground. (ii) Find the horizontal distance between O and the wall. [4]

OpenStudy (abmon98):

i solved i) but i am stuck on ii)

OpenStudy (shamim):

U know Horizontal component of initial velocity =v*cos theta

OpenStudy (shamim):

Vertical component of initial velocity=v*sin theta

OpenStudy (abmon98):

yes

OpenStudy (shamim):

Use for y component y=v*sin theta*t-0.5gt^2 Given y=4m

OpenStudy (abmon98):

shouldnt we take into accout the maximum height

OpenStudy (shamim):

Ya

OpenStudy (shamim):

x=v*cos theta*t=?

OpenStudy (shamim):

I think u know initial velocity frm i)

OpenStudy (abmon98):

for i) at max height v=0 m/s Vertical velocity vsinθ-gt=0 and horizontal displacement of vcosθt=19.2 t=19.2/vcosθ plug that back into the vertical velocity vector. vsinθ-g(19.2)/vcosθ=0 cosθ=4/5 and sinθ=3/5. solve for v, v=20m/s^2.

OpenStudy (abmon98):

sorry m/s

OpenStudy (shamim):

Ok

OpenStudy (abmon98):

20cosθt=19.2 t=19.2/16=1.2 seconds 20sinθt-5t^2=hmax hmax=7.2m

OpenStudy (abmon98):

ii) u=0m/s a=gm/s^2 s=4m v=vm/s use v^2=u^+2as v=4root5 m/s

OpenStudy (abmon98):

vertical displacement:7.2-vsinθt-1/2gt^2=4 (7.2-4)+4root5*3/5*t-5t^2=0

OpenStudy (abmon98):

t=1.499 seconds

OpenStudy (shamim):

Ya

OpenStudy (shamim):

Now x=v*cos theta*t=?

OpenStudy (abmon98):

when i checked out my answers the distance from O to the wall is incorrect but 4root 5 the speed is correct.

OpenStudy (shamim):

I hv little confusion!!!

OpenStudy (shamim):

U wrote Vertical displacement 7.2. Why?!!!

OpenStudy (shamim):

Can u use y=v*sin theta*t-0.5gt^2 Given y=4m v=20 t=?

OpenStudy (abmon98):

7.2 m is the maximum height

OpenStudy (shamim):

I c

OpenStudy (abmon98):

why is speed equal to 20m/s shouldnt it equal to 4root5

OpenStudy (shamim):

That wrong to use h=7.2-ut-0.5gt^2

OpenStudy (abmon98):

okay then have we chosen speed to be equal to 20m/s

OpenStudy (shamim):

U calculated the initial velocity of projection v=20m/s Right?

OpenStudy (abmon98):

yes thats the initial as it projects upwards till it reaches max height of which then velocity equals zero then velocity increases again

OpenStudy (abmon98):

yes i did.

OpenStudy (shamim):

But this is initial velocity of projection v=20m/s

OpenStudy (shamim):

Now can u tell me t=? Frm my equation

OpenStudy (abmon98):

2 seconds

OpenStudy (shamim):

Now use x=v*cos theta*t=?

OpenStudy (abmon98):

v=32m/s

OpenStudy (shamim):

v=32m/s?!!!

OpenStudy (shamim):

Did u get it???

OpenStudy (abmon98):

i still don't get why the intital velocity of projection should be the same after the particle reaches maximum height.

OpenStudy (abmon98):

At the instant when the stone hits the wall the horizontal component of the stone’s velocity is halved in magnitude and reversed in direction. The vertical component of the stone’s velocity does not change as a result of the stone hitting the wall. (iii) Find the distance from the wall of the point where the stone reaches the ground.

OpenStudy (shamim):

Actually velocity of projection is not equal to the velocity at highest point!!

OpenStudy (shamim):

U know any stone going upward and its velocity decreases

OpenStudy (shamim):

But my question is... why u r using the velocity of highest point?

OpenStudy (shamim):

Actually ur initial velocity of projection is v

OpenStudy (shamim):

X comonent of v is v*cos theta. And this component of velicity is responsible for displacement along x axis

OpenStudy (shamim):

Y component of v is v*sin theta. And this component is responsible for displacement along y axis

OpenStudy (shamim):

Our actual equation is h=ut-0.5gt^2 But we write this equation here as y=v*sin theta*t-0.5*g*t^2

OpenStudy (shamim):

I use u=v*sin theta Because of displacement along y axis

OpenStudy (shamim):

I wanna clear ur confusion!!!

OpenStudy (abmon98):

its crystal clear now thanks @shamim.

OpenStudy (shamim):

U r most welcome!!!!!